find the first five terms of the sequence in which a1=4 and an+1=3an+5, n is grater than or equal to 1.
i hope you can help
thanks a bunch!

Answer by stanbon(75887)

find the first five terms of the sequence in which a1=4 and an+1=3an+5, n is grater than or equal to 1.
--------------
a(1)=4
a(2)=3(4)+5=17
a(3)=3(17)+5=56
a(4)=3(56)+5=173
a(5)=I'll leave it to you.
Cheers,
Stan H.